CSePub Home Book Store Instructors Indopendent Authors About you new house The tistcomprycNers tee instaldon a d egaprnerd, b tw.chage you S426 00 year ใจเever The secand company changes 5751 00 tar installation, bud wil charge you s242 00 per yoa torever. Assurme that payments are af the END of the year Your per year, and you want to evalate each proposal What is the PV of the paid instualation? (Thies anser will be a NEGATIVE as you are spending cash for the syntem without receving any payofs) Answer Format: Curvency: Round to: 2 decimal places Enter Answer Here Submit Answe
Answer #1
Present value of first company = Installation cost + ( Yearly charges / Discount rate)=
0 + ( 426 / 6% )= $7100
Present value of second company = Installation cost + ( Yearly charges / Discount rate)=

751 + (242/6%) = 751 + 4033.33 = $4784.33

Second company should be sellected as lower present value of cost.